# Faster Random Walk-based Capacitance Extraction with Generalized Antithetic Sampling

### v2

@@ Abstract
Floating random walk-based capacitance extraction has emerged in recent years as a tried and true approach for extracting parasitic capacitance in very large scale integrated circuits. Being a Monte Carlo method, its performance is dependent on the variance of sampled quantities and variance reduction methods are crucial for the challenges posed by ever denser process technologies and layout-dependent effects. In this work, we present a novel, universal variance reduction method for floating random walk-based capacitance extraction, which is conceptually simple, highly efficient and provably reduces variance in all extractions, especially when layout-dependent effects are present. It is complementary to existing mathematical formulations for variance reduction and its performance gains are experienced on top of theirs. Numerical experiments demonstrate substantial such gains of up to 30% in number of walks necessary and even more in actual extraction times compared to the best previously proposed variance reduction approaches for the floating random-walk.
@@ IV-A Motivation and High-level idea
Note that the scheme of [5] is conceptually relevant to antithetics, however for selecting the initial (not antithetic) sample, the support of the pdf on the surface of the transition domain is changed, since they only sample on a fraction of the surface of the domain, not the whole surface. This is the reason why they need to subsequently weigh the samples.
@@
$\begin{split}p(\mathbf{x_{1}},...,\mathbf{x_{k}})&=\prod\limits_{i=1}^{k}q(\mathbf{r},\mathbf{r_{i}})\\ p(\mathbf{x_{k}}\rvert\mathbf{x_{1}}...,\mathbf{x_{k-1}})&=p(\mathbf{x_{k}})=q(\mathbf{r},\mathbf{r_{i}})\end{split}$ (8)
@@ V Numerical Experiments
  • SMS-2: This is the multiple-shooting estimator of [5] employing two points on each first transition domain selected symmetrically with respect to the VGS.
  • GAS-2: This is the GAS estimator proposed in this paper, employing two points on each first transition domain, selected according to Algorithm 1.
In both the multiple-shooting variants (SMS-2 and GAS-2) we employed 2 walks per first transition domain to better isolate the essential idea of each method. The main reason for this is that there is a tradeoff between taking more samples on the same first transition domain and forming more (different) first transition domains on the VGS, which also depends on the type of design to be extracted and is also alluded to in [5]. For example, we have experienced substantial performance gains with GAS-2 over IS+SS in virtually every design we have tried, and the same cannot be claimed for GAS-4,6,8,10 compared to GAS-2. A case where GAS does indeed experience substantial improvement with increasing number of antithetic pairs is, e.g., a parallel plate capacitor design. We consider such additional improvement with respect to GAS-2 to be highly design-dependent, warranting an ablation study on its own and therefore designate it as part of future work and beyond the scope of this paper. Additionally, the theoretical results of [5] are centered specifically on the case of shooting 2 walks from each first transition domain and GAS-2 has been selected for a fair comparison.
Note that it is entirely possible functionally to take $N$ antithetic pairs per first transition domain with Algorithm 1, instead of just one, and these will be by construction independent to one another. Then, the total sample covariance will be a sum of covariance terms of the form (13) and as such negative as a whole. Furthermore, GAS allows any multiple of 2 as number of samples on the surface of each first transition domain (e.g., 4, 6, 8, 10), as opposed to SMS that allows only powers of 2 (e.g., 4, 8, 16, 32) and thus GAS is more flexible in comparison.
@@
The termination criterion for net extraction in the experiments was always relative standard error, i.e., the ratio of sample standard deviation over sample mean, to be less than 0.5% (i.e., 0.005). Several independent extractions for each design case were ran to mitigate the inherent stochasticity of the FRW method. The batch size utilized throughout was 1000 to achieve better granularity over possible differences in walks. The LDE cases utilized the machine learning models for Green’s function data from [2] to facilitate extraction.
Besides wall times, another performance metric examined was the number of first transition domains formed until convergence, which we consider to be a more objective metric of variance reduction for these types of approaches, all other things being equal, because it directly measures how many samples are necessary to achieve the same level of accuracy by a Monte Carlo estimator. Our implementation of the variants differed exclusively in the way samples were selected on the surface of the first transition domains, all other infrastructure was identical, including Green’s function data through precomputation, machine learning models as well as covariance calculations for convergence (of course, IS+SS completely forfeits such calculations). Hence we do not expect hops to differ substantially among the variants and focus on total number of first transition domains. The latter has a straightforward interpretation in terms of walks as well: for IS+SS it is exactly the number of walks, whereas for SMS-2 and GAS-2 walks can be found by doubling the number of first transition domains.
@@
The main results of experiments for all FRW variants are shown in Table II. Number of first transition domains reported are the sum of first transition domains to extract all nets in the design. Total capacitance $C_{tot}$ reported is the sum of cross capacitances (or equivalently the sum of self-capacitances) for all nets in the design.
@@
It can be seen that both SMS-2 and GAS-2 are orders of magnitude faster than IS+SS for the large cases 3 and 4 while also being noticeably faster in the simpler ones (cases 1 and 2). Cases 3 and 4 in particular, quite emphatically affirm the favorable performance results for SMS-2 presented in [5] as it achieves large speedups of 7x and 3.5x respectively over IS+SS. This also provides evidence for the validity of our implementation of the SMS-2 scheme. Our proposed GAS-2 scheme follows suit with only a very slight edge over SMS-2.
On the other hand, there is a distinct advantage of our proposed GAS-2 scheme in cases 4-9 over SMS-2 ranging from 9.5% to 30% in terms of first transition domains necessary for convergence, further scrutinized in Table 3, which offers an explanation for the improvement brought forth by GAS-2. The basic shortcoming of SMS-2 is it has no way of guaranteeing that taking symmetric points with respect to the Gaussian surface will lead to a total negative sample covariance in (13). This could happen primarily in cases such as 5-9, where the Green’s function itself exhibits asymmetries, because ambient dielectrics are not necessarily stratified and thus do not profess xy-plane rotational invariance. It could also presumably happen in designs with stratified dielectrics, but now on the top and bottom sides of the Gaussian surface, because there is no symmetry along the z-axis.
@@
A small percentage of symmetric pairs of points on the same first transition domains do in fact have the same sign and produce positive products for SMS-2 as shown in Table 3. This table also includes the value of $F$ as extra information on the size of the examined designs. Note that the sample covariance estimator of (13) is known to exhibit highly non-linear behavior. This means that the first term contributing to $\Delta$ may easily end up being closer to zero (i.e., ”less negative”) than in GAS-2 and hurt overall variance reduction.
TABLE III: Improvement of GAS-2 over SMS-2

### v3

@@ Abstract
Floating random walk-based capacitance extraction has emerged in recent years as a tried and true approach for extracting parasitic capacitance in very large scale integrated circuits. Being a Monte Carlo method, its performance is dependent on the variance of sampled quantities and variance reduction methods are crucial for the challenges posed by ever denser process technologies and layout-dependent effects. In this work, we present a novel, universal variance reduction method for floating random walk-based capacitance extraction, which is conceptually simple, highly efficient and provably reduces variance in all extractions, especially when layout-dependent effects are present. It is complementary to existing mathematical formulations for variance reduction and its performance gains are experienced on top of theirs. Numerical experiments demonstrate substantial such gains of up to 50% in number of walks necessary as well as in actual extraction times compared to the best previously proposed variance reduction approaches for the floating random-walk.
@@ IV-A Motivation and High-level idea
Note that the scheme of [5] is conceptually relevant to antithetics, however for selecting the initial (not antithetic) sample, the pdf itself and its support on the surface of the transition domain is changed, since they only initially sample on a fraction of the surface of the domain, not the whole surface. This is the reason why they need to subsequently weigh their samples.
@@
$\begin{split}p(\mathbf{x_{1}},...,\mathbf{x_{k}})&=\prod\limits_{i=1}^{k}q(\mathbf{r},\mathbf{r_{i}})\\ p(\mathbf{x_{k}}\rvert\mathbf{x_{1}}...,\mathbf{x_{k-1}})&=p(\mathbf{x_{k}})=q(\mathbf{r},\mathbf{r_{k}})\end{split}$ (8)
@@ V Numerical Experiments
  • SMS-n: This is the multiple-shooting estimator of [5] employing n=2,4,8n={2,4,8} points on each first transition domain selected symmetrically with respect to the VGS.
  • GAS-n: This is the GAS estimator proposed in this paper, employing n=2,4,8n={2,4,8} points on each first transition domain, selected according to Algorithm 1.
As already mentioned, it is entirely possible functionally to take $N$ antithetic pairs per first transition domain with Algorithm 1, instead of just one, and these will be by construction independent to one another. Then, the total sample covariance will be a sum of covariance terms of the form (13) and as such negative as a whole. Furthermore, GAS allows any multiple of 2 as number of samples on the surface of each first transition domain (e.g., 4, 6, 8, 10), as opposed to SMS that allows only powers of 2 (e.g., 4, 8, 16, 32) and thus GAS is more flexible in comparison.
@@
The termination criterion for net extraction in the experiments was always the relative standard error, i.e., the ratio of sample standard deviation over sample mean, to be less than 0.5% (i.e., 0.005). Several independent extractions for each design case were ran to mitigate the inherent stochasticity of the FRW method. The batch size utilized throughout was 1000 to achieve better granularity over possible differences in walks. The LDE cases utilized the machine learning models for Green’s function data from [2] to facilitate extraction.
Besides wall times, another performance metric examined was the number of first transition domains formed until convergence, which we consider to be a more objective metric of variance reduction for these types of approaches, all other things being equal, because it directly measures how many samples are necessary to achieve the same level of accuracy by a Monte Carlo estimator. Apart from the way samples were selected on the surface of the first transition domains, all other infrastructure for our implementation of the variants was identical, including Green’s function data through precomputation and machine learning models. Hence we do not expect hops to differ substantially among the variants and focus on total number of first transition domains. The latter has a straightforward interpretation in terms of walks as well: for IS+SS it is exactly the number of walks, whereas for SMS-n and GAS-n walks it can be found by multiplying the number of first transition domains by $n$.
@@
TABLE III: Results
SMS-4 GAS-4 SMS-8 GAS-8
Case #first domains time(sec) #first domains time(sec) #first domains time(sec) #first domains time(sec)
1 181600 14.7 145600 13 103700 15.4 73400 11.5
2 450000 48.9 450000 48.8 225000 41.2 225000 40.9
3 2212300 869.2 1659700 635.8 1761100 1225.2 1290800 863.4
4 3035000 3383 2033800 1816.6 1637800 3394.4 1105900 1907.9
5 94700 222.8 71200 158.4 50100 193.7 35900 139.3
6 58400 397.2 39200 257.4 35100 436.7 20000 234
7 151400 203.3 123300 157.4 80300 174.3 61300 128
8 67800 259.8 41800 157.3 40300 268.6 21100 137.1
9 32900 97.8 19000 58.9 20500 102.7 10000 51.5
@@
The main results of experiments for all FRW variants are shown in Tables II and III. The number of first transition domains reported is the sum of first transition domains to extract all nets in the design. Total capacitance $C_{tot}$ reported is the sum of cross capacitances towards conductors of different nets for all nets in the design. Due to space considerations, $C_{tot}$ is only reported for SMS-2 and GAS-2, since extracted values do not substantially differ for the other variants.
For both the multiple-shooting variants (SMS-n and GAS-n) there is a tradeoff between taking more samples on the same first transition domain and forming more (different) first transition domains on the VGS, which also depends on the type of design to be extracted and is also alluded to in [5]. For example, we have experienced substantial performance gains with GAS-2 over IS+SS in virtually every design we have tried, and the same cannot be claimed for GAS-4,8 compared to GAS-2. The best wall times for the largest cases 3 and 4 are not achieved by the GAS-8 variant, but by the GAS-2 and GAS-4 variants respectively. For the SMS schemes, extraction time actually deteriorates when moving from SMS-2 to SMS-4 in cases 3 and 4 and when moving from SMS-4 to SMS-8 in cases 3, 4, 6, 8 and 9. It is hence not straightforward that starting as many walks as possible from each first transition domain is the best approach for every design. We consider additional improvement with respect to GAS-2 and SMS-2 to be highly design-dependent, warranting an ablation study of its own to -ideally- predict whether a given design can benefit from such a choice.
@@
It can be seen that both SMS-n and GAS-n are orders of magnitude faster than IS+SS for the large cases 3 and 4, while also being noticeably faster in the simpler ones (cases 1 and 2). Cases 3 and 4 in particular, quite emphatically affirm the favorable performance results for SMS-n presented in [5] as it achieves large speedups of 7x and 3.5x respectively over IS+SS. Our proposed GAS-2 scheme follows suit with only a very slight edge over SMS-2. For cases 1-3, GAS-2 and SMS-2 are practically equivalent.
On the other hand, there is a distinct advantage of our proposed GAS-n scheme in cases 4-9 over SMS-n, ranging from 9.5% to 50% in terms of first transition domains necessary for convergence, as well as in terms of wall times.
Thorough investigation of why the SMS schemes are lagging in performance is largely outside the scope of this paper. To begin with, SMS variance in [5] uses a different mathematical formulation, where negative correlation among the samples is an emerging, not foundational property. We only describe possible reasons viewed on the basis of covariance (12)-(13), while examining SMS-2 and GAS-2 in particular, to better isolate the essential idea of each method. Under the covariance view, a basic shortcoming of SMS-2 is it has no way of guaranteeing that taking symmetric points with respect to the Gaussian surface will lead to a negative sample covariance in (13). This could happen primarily in cases such as 5-9, where the Green’s function itself exhibits asymmetries, because ambient dielectrics are not necessarily stratified and thus do not profess xy-plane rotational invariance. It could also presumably happen in designs with stratified dielectrics, but now on the top and bottom sides of the Gaussian surface, because there is no symmetry along the z-axis. A second possible and related issue of SMS-2 is that each weight used in the capacitance estimator has an extra multiplicative factor necessary to make the final estimator unbiased. These factors are not fixed across first transition domains and, under the covariance view, could in principle increase final variance. GAS-2 has by construction a fixed weight factor of 0.5, according to (9).
@@
The above reasons are further scrutinized in Table V, where the next to last column reports the percentages of all multiplicative weight factors that differ from 0.5 by more than 0.05. This table also includes the value of $F$ as extra information on the size of the examined designs. A small percentage of symmetric pairs of points on the same first transition domains do in fact have the same sign and produce positive products for SMS-2 as shown in Table V. Note that the sample covariance estimator of (13) is known to exhibit highly non-linear behavior. This means that the first term contributing to $\Delta$ may easily end up being closer to zero (i.e., ”less negative”) than in GAS-2 and hurt overall variance reduction. Furthermore, it can indeed be the case that a significant percentage of weight factors is not close to 0.5 (although not 100%). This is not a full distributional analysis of the weight factors of course, however, it does show that there are both factors close to 0.5 and factors deviating from 0.5, which can, in principle, lead to additional variance of the final SMS-2 estimator. Note that in cases 1-3 where SMS-2 is competitive, either the percentage of positive correlation products is low, or the percentage of non-trivial weight factors is low, or both.
The above comments indicate that the SMS schemes could be affected by factors to which the GAS schemes are agnostic, without compromising efficiency.
TABLE IV: Improvement of GAS-2 over SMS-2